# Minimal codes from hypersurfaces in even characteristic

### v3

@@ Abstract
The setting of projective systems can be used to study the parameters of a projective linear code $\mathscr{C}$. This can be done by considering the intersections of the point set $\Omega$ defined by the columns of a generating matrix for $\mathscr{C}$ with the hyperplanes of a projective space. In particular, $\mathscr{C}$ is minimal if $\Omega$ is cutting, i.e., every hyperplane is spanned by its intersection with $\Omega$. Minimal linear codes have important applications for secret sharing schemes and secure two-party computation.
@@
In the last part of the paper we consider an extension of the notion of being cutting with respect to subspaces other than hyperplanes and introduce the definition of cutting gap in order to characterize and measure what happens when this property is not satisfied. Finally, we then apply these notions to Hermitian codes and to the codes related to $\mathscr{V}_{\varepsilon}^{r}$ discussed before.
@@ 1 Introduction
Given a non-degenerate $q$-ary linear $[n,r,d]$-code $\mathscr{C}$, the projective system of $\mathscr{C}$ is the (multi)set $\Omega$ of points of $\mathrm{PG}(r-1,q)$ determined by the columns of any generator matrix of $\mathscr{C}$. Clearly, $\Omega$ is not uniquely determined, but it can be easily seen that it is unique up to projectivities. Conversely, given any (multi)set $\Omega$ of points in $\mathrm{PG}(r-1,q)$ it is possible to construct a code $\mathscr{C}(\Omega)$ whose generator matrix has as columns the coordinates of the points in $\Omega$. The code $\mathscr{C}$ is also not uniquely determined, but it turns out to be unique up to (monomial) code equivalence.
@@
An $[n,r,d]$-linear code $\mathscr{C}$ is minimal if all of its codewords are minimal that is are identified, up to a non-zero scalar multiple, by their support. The projective system $\Omega$ arising from a minimal code is a cutting set in $\mathrm{PG}(r-1,q)$, that is to say any hyperplane $\Pi$ of $\mathrm{PG}(r-1,q)$ is spanned by its intersection with $\Omega$. Minimal codes are relevant both for their geometric properties and possible applications to secret-sharing schemes and secure two-party computation; see [17, 10].
A secret sharing scheme is a cryptographic procedure allowing to handle the access to a secret by a group of people, the participants. Every participant has a share of the secret and the whole secret can be retrieved only when a suitable group of participants puts its shares together. Such a “suitable group of participants” is called access set, while the collection containing all access sets is called access structure for the given scheme. We also say that an access set is minimal, if removing from it any participant, it is not possible to find out the secret anymore. Massey, in [17], showed how linear error-correcting codes can be used to build secret sharing schemes showing the link between the access structures and the minimal words of the dual code.
In particular, the idea is that, taken a generator matrix $G$ for an $[n,r,d]$-code $\mathscr{C}$, a trusted party considers a vector $\boldsymbol{u}$ and computes the product $\boldsymbol{u}G=(t_{0},t_{1},\ldots,t_{n-1})$; the first entry of this vector is the secret $s$, that is $t_{0}={\boldsymbol{u}}g_{0}=s$. It is easily seen that a set of shares $\{t_{i_{1}},t_{i_{2}},\ldots,t_{i_{m}}\}$ determines the secret if and only if $g_{0}$ is a linear combination of $g_{i_{1}},\ldots,g_{i_{m}}$
A subset of shares $\{t_{i_{1}},t_{i_{2}},\ldots,t_{i_{m}}\}$ can reconstruct the secret if and only if there is a codeword $\boldsymbol{c}=(1,0,0,\ldots,c_{i_{1}},0,\ldots,c_{i_{m}},0,\ldots,0)$ in the dual code, such that $c_{i_{j}}\neq 0$ for some $j$. A minimal codeword then gives rise to a minimal access set. It is quite difficult to find out all minimal words in an arbitrary linear code and this difficulty translates in the difficulty in finding good access structures. Such structures are good since they are democratic, so every participant belongs to the same number of minimal access sets.
In the present paper we construct codes with few weights arising from hypersurfaces related to quasi-Hermitian varieties in even characteristic that in some cases are minimal and in others turn out to be $3$-weight codes. Linear codes with few weights also have applications to authentication codes and association schemes; see [8]. Furthermore, the study of the weights is important, since they measure the efficiency of the code and their knowledge is useful for decoding.
In the last part of the paper we consider an extension of the notion of being cutting with respect to subspaces other than hyperplanes, see [11], and introduce the definition of cutting gap in order to characterize and measure what happens when this property is not satisfied.
The notion of being cutting for a set is interesting also from a geometric point of view, as it is related to that of blocking sets, see [7]. In the last part of the paper we wish to consider extensions of the notion of being cutting to subspaces of codimension greater than $1$. In general, being cutting with respect to arbitrary subspaces turns out to be a very strong requirement. In order to better manage the geometry we will introduce the notion of cutting gap to describe how far a set $\Omega$ is from being cutting in codimension greater than one. This has a twofold aim: on one side to investigate the geometry of $\Omega$ as set of points embedded in $\mathrm{PG}(r-1,q)$ at a “finer” scale than what is provided by the hyperplane sections; on the other to start a project to look for analogues of being minimal for a code with respect to higher Hamming weights.
The paper is organized as follows. In Section 2 we recall basic results on projective codes and quasi-Hermitian varieties; in particular we consider Buekenhout-Tits (BT) quasi-Hermitian varieties which are proven to exist in any $\mathrm{PG}(r,q^{2})$ with $r>1$, $q=2^{e}$ and $e$ odd.
Next, in Section 3 we further investigate some hypersurfaces $\mathscr{V}^{r}_{\varepsilon}$ related to BT quasi-Hermitian varieties and prove that they are cutting (so the related codes are minimal) if and only if $r=3$ and $e\equiv 3\pmod{4}$ or $r\geq 4$. For $r=3$ we compute the weights of the code $\mathscr{C}_{\varepsilon}^{3}:=\mathscr{C}(\mathscr{V}^{3}_{\varepsilon})$ and for $r\geq 3$ provide a new family of codes with just $3$ weights.
Finally, in Section 4 we formally introduce the notion of cutting gap and propose some problems related to it. Next, we apply these notions to Hermitian codes. We also investigate the intersection with planes and lines of ${\mathscr{V}}^{4}_{\varepsilon}$ in $\mathrm{PG}(4,q^{2})$. In this latter case, we analyze also some properties of Fermat curves. As a byproduct, we provide the sequence of cutting gaps of ${\mathscr{V}}^{4}_{\varepsilon}$ and the complete list of the weights of the code $\mathscr{C}_{\varepsilon}^{4}:=\mathscr{C}(\mathscr{V}^{4}_{\varepsilon})$.
@@ 2.1 Coding Theory
Let $\Omega$ be a set of $N$ points of $\mathrm{PG}(r,q)$ such that $\langle\Omega\rangle=\mathrm{PG}(r,q)$. For each $i\in\{1,\ldots,N\}$, let $\boldsymbol{\textnormal{v}}_{i}\in\mathbb{F}_{q}^{r+1}$, be a fixed representative of a point $P_{i}=\langle\textbf{{v}}_{i}\rangle\in\Omega$ and denote by $\mathscr{C}(\Omega)$ the projective linear code having generator matrix whose columns are the vectors $\textbf{{v}}_{i}$. Then $\mathscr{C}(\Omega)$ is a minimal code if and only if for any hyperplane $\Pi$ of $\mathrm{PG}(r,q)$,
@@
Clearly, $d_{1}(\mathscr{C})$ corresponds to the usual minimum Hamming distance $d$ of $\mathscr{C}$. More in general, it can be seen that for $11\leq k\leq r-$ 1 the $k$–th Hamming weight of a projective code $\mathscr{C}=\mathscr{C}(\Omega)$ is given by
@@ 2.2 Quasi-Hermitian varieties
Quasi-Hermitian varieties in $\mathrm{PG}(r,q^{2})$ are point sets having the same intersection numbers with respect to hyperplanes as a non-singular Hermitian variety $\mathscr{H}(r,q^{2})$ of $\mathrm{PG}(r,q^{2})$. Notice that a point set $\mathscr{S}$ of $\mathrm{PG}(r,q^{2})$, $r>2$, having the same intersection numbers with respect to hyperplanes as a non‐singular Hermitian variety has also the same number of points as $\mathscr{H}(r,q^{2})$ (see [18]). They are an extension to higher dimensions of the well-known notion of unital in the projective plane.
More in detail, the cardinality of a quasi-Hermitian variety of $\mathrm{PG}(r,q^{2})$ is
$|\mathscr{H}(r,q^{2})|=\frac{(q^{r+1}+(-1)^{r})(q^{r}-(-1)^{r})}{(q^{2}-1)},$
@@
$|\Pi_{0}\mathscr{H}(r-2,q^{2})|=\frac{(q^{r}+(-1)^{r-1})(q^{r-1}-(-1)^{r-1})}{q^{2}-1}+(-1)^{r-1}q^{r-1}.$
Here $\Pi_{i}$ denotes an $i$-dimensional space of $\mathrm{PG}(r,q^{2})$ and $\Pi_{0}\mathscr{H}(r-2,q^{2})$ is a cone, the join of the vertex $\Pi_{0}$ to a non-singular Hermitian variety $\mathscr{H}(r-2,q^{2})$ in a projective subspace $\Pi_{r-2}$ which does not contain $\Pi_{0}$. A non-singular Hermitian variety of $\mathrm{PG}(r,q^{2})$ is, by definition, a quasi-Hermitian variety: the classical quasi-Hermitian variety.
@@ 3.1 The variety $\mathscr{V}_{\varepsilon}^{3}$
Here we wish to study the projective linear code $\mathscr{C}^{3}_{\varepsilon}=\mathscr{C}(\mathscr{V}^{3}_{\varepsilon})$ induced by the projective system of the $\mathbb{F}_{q^{2}}$-rational points of $\mathscr{V}^{3}_{\varepsilon}\subset\mathrm{PG}(3,q^{2})$.
Firstly, we investigate the variety at infinity $\mathscr{V}_{\varepsilon,\infty}^{3}:=\mathscr{V}_{\varepsilon}^{3}\cap\Pi_{\infty}$.
@@
$|\mathscr{V}^{3}_{\varepsilon}|=\begin{cases}q^{5}+q^{2}+1&if\quad e\equiv 1\pmod{4}\\ q^{5}+3q^{2}+1&if\quad e\equiv 3\pmod{4}.\end{cases}$ (5)
@@
The code $\mathscr{C}_{\varepsilon}^{3}=\mathscr{C}(\mathscr{V}^{3}_{\varepsilon})$ is a $[|\mathscr{V}_{\varepsilon}^{3}|,4]$-linear code with minimum distance $d=|\mathscr{V}_{\varepsilon}^{3}|-q^{3}-q^{2}-1$. Moreover, its $m$-th higher minimum weights $d_{m}$ with $m\in\{1,2,3\}$ are:
@@ 3.2 The variety $\mathscr{V}_{\varepsilon}^{r}$, $r\geq 4$
Then, this variety is a cone through the point $P_{\infty}(0,0,\ldots,1)$ and a Fermat hypersurface $\mathscr{F}^{r-2}_{n}$, $n=2^{\frac{e-1}{2}}+1$, contained in the $(r-2)$-dimensional subspace $\Pi_{r-2}:X_{0}=X_{r}=0$. By Lemma 3.7, this variety spans $\Pi_{r-2}$ for $e\geq 3$. So, since $P_{\infty}$ belongs to $\mathscr{V}^{r}_{\varepsilon,\infty}\setminus\Pi_{r-2}$, we have
@@
    $\begin{split}(q^{2}-1)&\left(|\mathscr{H}^{r}_{\varepsilon}\cap\mathrm{AG}(r,q^{2})\cap\Pi|-\theta_{q^{2}}(r-2)\right)\\ &=(q^{2}-1)q^{r-2}(q^{r-1}-(-1)^{r-1})-(q^{2(r-1)}-1)\\ &=\begin{cases}(q^{r-1}+1)(q^{r}-q^{r-2}-q^{r-1}+1)&\,r\textnormal{ even }\\ (q^{r-1}-1)(q^{r}-q^{r-2}-q^{r-1}-1)&\,r\textnormal{ odd }.\end{cases}\end{split}$ (14)
@@
By Theorem 3.10, in order to get the length and the weights of the projective linear code $\mathscr{C}_{\varepsilon}^{r}$, $r\geq 4$, it is necessary to determine the size of the Fermat hypersurface $\mathscr{F}_{n}^{r-2}$, $r\geq 4$ and $n=2^{\frac{e-1}{2}}+1$, lying in the $(r-2)$-dimensional subspace of $\mathrm{PG}(r,q^{2})$ of equations $X_{0}=X_{r}=0$. Thus, we prove the following.
@@
    In this case, the plane curve $\mathscr{F}^{2}_{3}$ turns out to be maximal by Theorem [13, Theorem 10.65]. So the result follows from the Hasse-Weil bound.
Theorem 3.10.
$|\mathscr{V}^{r}_{\varepsilon}|=q^{2r-1}+|\mathscr{F}^{r-2}_{n}|q^{2}+1,$
where $n=2^{\frac{e-1}{2}}+1$.
Proof.
The result follows observing that $\mathscr{V}^{r}_{\varepsilon}\cap\Pi_{\infty}$ is a cone with vertex the point $P_{\infty}(0,0,\ldots,1)$ over a Fermat hypersurface of the $(r-2)$-dimensional subspace with equations $X_{0}=X_{r}=0$. ∎
@@
Now, consider the multiset $\overline{\mathscr{V}_{\varepsilon}^{r}}$ in $\mathrm{PG}(r,q^{2})$, $r\geq 3$, that consists of the points in $(\mathscr{V}_{\varepsilon}^{r}\setminus\Pi_{\infty})$ plus the point $P_{\infty}$ assigned multiplicity equal to $j\geq 1$. Using an approach similar to that used in [3] we get the following result.
@@
The linear code associated with $\overline{\mathscr{V}_{\varepsilon}^{r}}$ is a $[q^{2r-1}+j,r+1]_{q^{2}}$ code with weights:

### v4

@@ Abstract
The setting of projective systems can be used to study the parameters of a projective linear code $\mathscr{C}$. This can be done by considering the intersections of the point set $\Omega$ defined by the columns of a generating matrix for $\mathscr{C}$ with the hyperplanes of a projective space. In particular, $\mathscr{C}$ is minimal if $\Omega$ is cutting set, i.e., every hyperplane is spanned by its intersection with $\Omega$. Minimal linear codes have important applications for secret sharing schemes and secure two-party computation.
@@
In the last part of the paper, we consider an extension of the notion of being cutting with respect to subspaces other than hyperplanes and introduce the definition of a cutting gap in order to characterize and measure what happens when this property is not satisfied. Finally, we apply these concepts to both Hermitian varieties and the hypersurfaces $\mathscr{V}_{\varepsilon}^{r}$.
Keywords: Algebraic variety, Quasi-Hermitian variety, linear code.
MSC Classification: 94B05, 51A05, 51E20.
@@ 1 Introduction
Linear codes derived from projective geometry have been extensively investigated due to the rich interplay between algebraic and geometric structures. The approach of constructing generator matrices for codes by considering the coordinates of point-sets in projective spaces has been pioneered by [19] and [8] and it has proved to be extremely fruitful ever since. In particular, given a non-degenerate $q$-ary linear $[n,r,d]$-code $\mathscr{C}$, the projective system associated with $\mathscr{C}$ is defined as the (multi)set $\Omega$ of points in $\mathrm{PG}(r-1,q)$ corresponding to the columns of a generator matrix of $\mathscr{C}$. Conversely, any such multiset $\Omega$ yields a code $\mathscr{C}(\Omega)$ by taking the coordinates of the points in $\Omega$ as the columns of a generator matrix. The resulting code $\mathscr{C}(\Omega)$ is unique up to monomial equivalence and is referred to as the code generated or induced by $\Omega$.
This geometric approach is particularly effective when $\Omega$ arises from algebraic or geometric structures, as it often leads to codes with desirable features, such as minimality or a small number of distinct weights relative to the code length.
@@
A linear $[n,r,d]$-code is said to be minimal if every non-zero codeword is minimal, i.e., uniquely determined (up to a non-zero scalar multiple) by its support. Geometrically, this corresponds to the condition that the associated projective system $\Omega$ is a cutting set, meaning that every hyperplane of $\mathrm{PG}(r-1,q)$ is spanned by its intersection with $\Omega$. Minimal codes are of interest not only for their geometric significance but also for their applications in cryptography, including secret-sharing schemes and secure two-party computation, see [20, 12].
In this paper, we construct new families of linear codes with few weights, derived from hypersurfaces associated with quasi-Hermitian varieties in even characteristic. These constructions are noteworthy for their geometric novelty. Specifically, the hypersurfaces considered are linked to Buekenhout–Tits quasi-Hermitian varieties (or BT quasi-Hermitian varieties, for short), which exist in every $\mathrm{PG}(r-1,q^{2})$ with $r>2$ and $q=2^{e}$, where $e\geq 3$ is odd. The structure of these varieties enables the exploration of new geometric configurations in finite projective spaces.
In certain cases, the associated codes are minimal; in others, they exhibit exactly three distinct weights. Codes with few weights are particularly valuable in the design of authentication codes and association schemes (see [9]), and their weight distribution plays a crucial role in evaluating decoding performance and efficiency.
In the final part of the paper, we introduce a new concept, the cutting gap, to quantify the extent to which a set $\Omega$ fails to satisfy the cutting property (i.e. that a subspace is generated by its intersection with $\Omega$). This notion aims to quantify how much a given projective system $\Omega$ deviates from being a strong blocking set (see [13]) and also aims to open the way to extending the concept of minimality to higher Hamming weights.
The paper is organized as follows. In Section 2, we recall foundational results on projective codes and quasi-Hermitian varieties, with particular emphasis on BT quasi-Hermitian varieties. Section 3 is devoted to the analysis of hypersurfaces $\mathscr{V}^{r}_{\varepsilon}$ associated with these varieties, and we determine the conditions under which they yield minimal codes. For $r=3$, we compute the nonzero weights of $\mathscr{C}_{\varepsilon}^{3}:=\mathscr{C}(\mathscr{V}^{3}_{\varepsilon})$, while for $r\geq 3$, we construct a new family of three-weight codes.
In Section 4, we formally define the concept of cutting gap and explore its implications. We apply this framework to Hermitian varieties and study the intersections of $\mathscr{V}^{4}_{\varepsilon}$ with planes and lines in $\mathrm{PG}(4,q^{2})$, including an analysis of Fermat curves. As a result, we provide the complete sequence of cutting gaps for $\mathscr{V}^{4}_{\varepsilon}$ and the full list of weights for the code $\mathscr{C}_{\varepsilon}^{4}:=\mathscr{C}(\mathscr{V}^{4}_{\varepsilon})$. In both cases, $r=3$ or $r=4$, the limit
$\lim_{q\to\infty}\frac{N(q)}{d(q)}=1,$
where $N(q)$ denotes the code length and $d(q)$ the minimum distance. This implies that the asymptotic minimum distance of the constructed codes is relatively good. In the following table, we summarize all the results obtained concerning the code $\mathscr{C}_{\varepsilon}^{r}$.
Code Result
  $\mathscr{C}^{3}_{\varepsilon}$ Theorem 3.2: length
Theorem 3.3: weights for $e\equiv 3\pmod{4}$
Theorem 3.4: weights for $e\equiv 1\ \pmod{4}$
Theorem 3.5: minimality for $e\equiv 3\pmod{4}$
Corollary 3.6: $m$-th higher minimum weights
$\mathscr{C}^{r}_{\varepsilon},r\geq 4$ Theorem 3.12 : minimality for $e$ odd
$\mathscr{C}(\overline{\mathscr{V}_{\varepsilon}^{r}})$ Theorem 3.13: weights
$\mathscr{C}^{4}_{\varepsilon}$ Proposition 4.9: minimality, length and weights
@@ 2.1 Coding Theory
Let $\Omega$ be a set of $N$ points of $\mathrm{PG}(r-1,q)$ such that $\langle\Omega\rangle=\mathrm{PG}(r-1,q)$. For each $i\in\{1,\ldots,N\}$, let $\boldsymbol{\textnormal{v}}_{i}\in\mathbb{F}_{q}^{r}$, be a fixed representative of a point $P_{i}=\langle\textbf{{v}}_{i}\rangle\in\Omega$ and denote by $\mathscr{C}(\Omega)$ the projective linear code having generator matrix whose columns are the vectors $\textbf{{v}}_{i}$. Then $\mathscr{C}(\Omega)$ is a minimal code if and only if for any hyperplane $\Pi$ of $\mathrm{PG}(r-1,q)$,
@@
Clearly, $d_{1}(\mathscr{C})$ corresponds to the usual minimum Hamming distance $d$ of $\mathscr{C}$. More in general, it can be seen that for $1\leq k\leq r-1$ the kk–th Hamming weight of a projective code $\mathscr{C}=\mathscr{C}(\Omega)$ is given by
@@ 2.2 Quasi-Hermitian varieties
Quasi-Hermitian varieties in $\mathrm{PG}(r,q^{2})$ are point sets having the same intersection numbers with respect to hyperplanes as a non-singular Hermitian variety $\mathscr{H}(r,q^{2})$ of $\mathrm{PG}(r,q^{2})$. Notice that a point set $\mathscr{S}$ in $\mathrm{PG}(r,q^{2})$, with $r>2$, having the same intersection numbers with respect to hyperplanes as a non-singular Hermitian variety also has the same number of points as $\mathscr{H}(r,q^{2})$ (see [21]).
These varieties can be seen as a natural extension to higher dimensions of the notion of a unital in the projective plane.
A unital embedded in $\mathrm{PG}(2,q^{2})$ is a set of $q^{3}+1$ points such that each line intersects it in either $1$ or $q+1$ points. The Hermitian curve $\mathscr{H}(2,q^{2})$ is a unital, known as the classical unital.
More precisely, the cardinality of a quasi-Hermitian variety in $\mathrm{PG}(r,q^{2})$ is
$|\mathscr{H}(r,q^{2})|=\frac{(q^{r+1}+(-1)^{r})(q^{r}-(-1)^{r})}{q^{2}-1},$
@@
$|\Pi_{0}\cap\mathscr{H}(r-2,q^{2})|=\frac{(q^{r}+(-1)^{r-1})(q^{r-1}-(-1)^{r-1})}{q^{2}-1}+(-1)^{r-1}q^{r-1},$
where $\Pi_{0}$ denotes a hyperplane tangent to the variety.
A non-singular Hermitian variety of $\mathrm{PG}(r,q^{2})$ is, by definition, a quasi-Hermitian variety: the classical quasi-Hermitian variety.
@@ 3 A family of minimal linear $q$-ary codes, $q$ even
Let $H$ be a quasi-Hermitian variety of $\mathrm{PG}(r,q^{2})$. As shown in [2], the linear code $\mathscr{C}(H)$ generated by $H$ is minimal. This was established by applying Theorem 2.2, since the sufficient condition for minimality given by Ashikhmin and Barg (1) could not be used.
Consider the BT quasi-Hermitian variety $\mathscr{H}_{\varepsilon}^{r}=(\mathscr{V}_{\varepsilon}^{r}\setminus\Pi_{\infty})\cup\mathscr{\mathscr{H}_{\varepsilon,\infty}^{r}}$. We ask whether $\mathscr{V}_{\varepsilon}^{r}$ induces minimal codes.
@@ 3.1 The variety $\mathscr{V}_{\varepsilon}^{3}$
We first study the projective linear code $\mathscr{C}^{3}_{\varepsilon}=\mathscr{C}(\mathscr{V}^{3}_{\varepsilon})$. As an initial step, we investigate the variety at infinity, defined as $\mathscr{V}^{3}_{\varepsilon,\infty}:=\mathscr{V}^{3}_{\varepsilon}\cap\Pi_{\infty}$.
@@
[…]
@@
The code $\mathscr{C}_{\varepsilon}^{3}=\mathscr{C}(\mathscr{V}^{3}_{\varepsilon})$ is a $|\mathscr{V}_{\varepsilon}^{3}|,4]-$linear code with minimum distance $d=|\mathscr{V}_{\varepsilon}^{3}|-q^{3}-q^{2}-1$. Moreover, its $m$-th higher minimum weights $d_{m}$ with $m\in\{1,2,3\}$ are:
@@
Remark 3.7.
We stress that in [11], the authors provided a general method to transform a minimal code satisfying the Ashikhmin–Barg condition into a minimal code that violates it.
@@ 3.2 The variety $\mathscr{V}_{\varepsilon}^{r}$, $r\geq 4$
Then, this variety is a cone through the point $P_{\infty}(0,0,\ldots,1)$ and a Fermat hypersurface $\mathscr{F}^{r-2}_{n}$, $n=2^{\frac{e-1}{2}}+1$, contained in the (r2)(r-2)-dimensional subspace $\Pi_{r-2}:X_{0}=X_{r}=0$. By Lemma 3.8, this variety spans $\Pi_{r-2}$ for $e\geq 3$. So, since $P_{\infty}$ belongs to $\mathscr{V}^{r}_{\varepsilon,\infty}\setminus\Pi_{r-2}$, we have
@@
[…]
@@
Theorem 3.10.
$|\mathscr{V}^{r}_{\varepsilon}|=q^{2r-1}+|\mathscr{F}^{r-2}_{n}|q^{2}+1,$
where $n=2^{\frac{e-1}{2}}+1$.
Proof.
The result follows observing that $\mathscr{V}^{r}_{\varepsilon}\cap\Pi_{\infty}$ is a cone with vertex the point $P_{\infty}(0,0,\ldots,1)$ over a Fermat hypersurface of the (r2)(r-2)-dimensional subspace with equations $X_{0}=X_{r}=0$. ∎
By Theorem 3.10, in order to get the length and the weights of the projective linear code $\mathscr{C}_{\varepsilon}^{r}$, $r\geq 4$, it is necessary to determine the size of the Fermat hypersurface $\mathscr{F}_{n}^{r-2}$, $r\geq 4$ and $n=2^{\frac{e-1}{2}}+1$, lying in the (r2)(r-2)-dimensional subspace of $\mathrm{PG}(r,q^{2})$ of equations $X_{0}=X_{r}=0$. Thus, we prove the following.
@@
    In this case, the plane curve $\mathscr{F}^{2}_{3}$ turns out to be maximal by Theorem [15, Theorem 10.65]. So the result follows from the Hasse-Weil bound.
@@
Now, consider the multiset $\overline{\mathscr{V}_{\varepsilon}^{r}}$ in $\mathrm{PG}(r,q^{2})$, $r\geq 3$, that consists of the points in $(\mathscr{V}_{\varepsilon}^{r}\setminus\Pi_{\infty})$ plus the point $P_{\infty}$, which is assigned multiplicity $j\geq 1$. Using an approach similar to that used in [3] we get the following result.
@@
The linear code associated with $\overline{\mathscr{V}_{\varepsilon}^{r}}$ is a $q^{2r-1}+j,r+1]_{q^{2}}$ code with weights: